# EpiCache: Episodic KV Cache Management for Long Conversational Question Answering

### v1

@@ 4.3 Ablation Study
  • Alternative design: replacing episodic caches with a RAG-like approach that directly inputs clustered conversation segments with the query.
@@ 4.4 Efficiency Analysis
Figure 7(a) reports decoding latency breakdown and peak GPU memory usage, comparing full KV with EpiCache under cache budgets of 2K-8K. Latency is evaluated per turn, consisting of answer decoding [^4] For fair comparison, decoding latency is measured by generating up to 10 tokens per turn. and, for EpiCache, query embedding, centroid matching, and KV cache retrieval. With fewer KV entries stored, EpiCache reduces decoding latency by up to $2.4\times$ and peak GPU memory by $3.5\times$ compared to full KV. These results demonstrate that EpiCache achieves both faster decoding and substantially lower memory usage, as block prefill bounds peak memory to the fixed cache budget throughout prefill and decoding.
Episode switching analysis in Figure 7(b) shows that retrieval overhead in EpiCache arises only when the matched episode changes between turns. We measure this by clustering each conversation, embedding subsequent turns, and tracking how often the assigned episode switches. The cumulative switch counts grow sublinearly compared to the Worst Case line (episode switched every turn), indicating that consecutive turns remain within the same episode. This conversational persistence amortizes retrieval cost, and when incorporated into the per-turn latency breakdown of Figure 7(a), the effective overhead remains under 5% while maintaining robust LongConvQA accuracy. Detailed measurements information are provided in Section D.1.
@@
Figure 7: Efficiency Analysis in Multi-Turn Conversation: (a) Per-turn decoding latency and peak GPU memory for full KV (100K) and EpiCache (E=4E{=}4) with LLaMA-3.2-3B. Query Embed and Match: query encoding and centroid matching, KVs Retrieve: loading episodic cache from CPU to GPU memory. (b) Cumulative episode switches in Realtalk with E=4E{=}4, showcasing how often episodes change across turns.
@@ 6 Conclusion
EpiCache is a framework that combines block-wise prefill with episodic clustering and sensitivity-aware budget allocation to preserve topic-relevant context under fixed memory. Across multiple LongConvQA benchmarks, EpiCache substantially outperforms existing compression methods, showing that efficient multi-turn interaction is feasible even under strict resource constraints and marking a practical step toward memory-efficient conversational AI.
@@ A.1 Dataset
We evaluate EpiCache on three LongConvQA benchmarks: Realtalk (Lee et al., 2025), LoCoMo (Maharana et al., 2024), and LongMemEval (Wu et al., 2025a). Three benchmarks follow the LongConvQA formulation in Section 2.1, where a long conversational history $\mathcal{H}$ is provided and the model is required to answer a sequence of queries $\mathcal{Q}={q_{1},\dots,q_{N_{q}}}$ grounded in dialogue history. This formulation is intended to assess the answer accuracy of LLMs in a multi-turn conversation.
@@
LoCoMo (Maharana et al., 2024) is a benchmark of long-term conversations, created through a human-machine pipeline where LLM-based agents generate dialogues grounded in distinct personas and temporal event graphs, and human annotators refine them for long-range consistency. The dataset consists of 10 conversations, each spanning up to 35 sessions with around 300 turns. The QA benchmark is divided into five subtasks: (i) Single-hop, (ii) Multi-hop, (iii) Temporal reasoning, (iv) Open-domain knowledge, and (v) Adversarial. Evaluation is conducted with open-ended genetation and report F1 score.
We exclude the adversarial subtask for the following reason. This task evaluates whether a model can detect unanswerable questions by choosing between a plausible but incorrect answer and a “no such information” response. Under KV cache compression, however, models often over-predict the latter, which spuriously inflates adversarial scores. For instance, on LLaMA-3.2-3B the score is only 12.11 with full KV, but rises to 49.78 under 4K KVzip compression. Unlike the consistent degradation seen in other subtasks (e.g., temporal or multi-hop reasoning), this inflation does not indicate real gains. Since open-source models already struggle on this task (Maharana et al., 2024), including it would be misleading. We therefore omit adversarial results from our main evaluation and defer a more careful study of unanswerability detection under compression to future work.
@@ A.2 KV Cache Compression Baseline Setup
Following Xiao et al. (2024), we retain a fixed number of sink and recent tokens throughout block prefill. Specifically, we fix the number of sink tokens to 128 for all models, while the remaining budget M128M-128 is assigned to the most recent tokens.
@@ A.3 EpiCache Setup
For segment construction, we set the embedding window size $w_{\text{embed}}$ to 4, selected from 2,4,8{2,4,8}. We observed that performance differences among these values were minor, making $w_{\text{embed}}{=}4$ a reasonable default. To cluster segments into episodes, we apply the standard K-means algorithm with k-means++ initialization (Arthur & Vassilvitskii, 2007). This offline segmentation and clustering stage completes within one minute, incurring negligible overhead.
For sensitivity-aware budget allocation, we estimate per-layer weights using a single randomly sampled conversation from the BookSum (Kryściński et al., 2022) dataset. By performing two forward passes—one with the full causal mask and one with the block-prefill mask—we measure the layer-wise deviations and compute allocation weights. Because only one sample is used, the overhead of this calibration step is negligible.
@@
For the patched prompt, we use the medoid segment of size 4, selected from 4,8,12{4,8,12}. Ablation studies show that different segment sizes yield only marginal performance differences. After constructing episodic KV caches, we offload them to offline memory (e.g., CPU) to minimize GPU memory usage. During online decoding, episodic caches are retrieved from offline memory. This design enables constrained GPU memory usage while keeping retrieval overhead manageable.
@@ Appendix B Alternative Design Exploration
A representative alternative design is a RAG-like approach, where conversation segments are clustered and the query is used to select the most relevant clusters, which are then directly fed into the LLM to form a KV cache under a fixed budget. The key difference from block prefill eviction is that block prefill spans the entire dialogue and selects tokens in a context-holistic manner, whereas the RAG-like method isolates only the chosen clusters when constructing the KV cache. As shown in Figure A1 (a), this leads to substantially worse performance—not only below EpiCache but also below block prefill-based compression methods such as KVzip.
This result further suggests that clustering alone is insufficient for directly constructing episodic memory inputs: conversation histories segmented only by clustering remain under-contextualized. Developing more effective clustering or episode partitioning strategies tailored to retrieval thus remains an interesting direction for future work.
@@
Figure A1 (b) varies the block size $M_{\text{block}}$ used during block prefill. Accuracy differences remain minor across settings. In practice, smaller block sizes reduce peak memory but require more iterations to span the full context, whereas larger blocks incur higher memory during prefill but process the history more efficiently. This reflects a memory-throughput trade-off, and block size can be tuned according to deployment constraints.
@@
Figures Figure A1 (c-f) examine clustering-related hyper-parameters. (i) Figure A1 (c) sweeps the segmentation window size $w_{\text{emb}}$ for utterance grouping, showing little sensitivity and indicating robustness to segmentation granularity. (ii) Figure A1 (d) varies the encoder for clustering and query embedding. The LLM’s own embedding layer yields weaker performance, while lightweight dedicated encoders (MiniLM-L6-v2, 0.06B; Qwen3-0.6B) deliver strong gains over InfiniPot. Larger encoders (Qwen3-4B) add only marginal improvements, suggesting that small encoders suffice for EpiCache. (iii) Figure A1 (e) sweeps the number of episodes $E$, where larger $E$ improves segmentation and accuracy—especially under tight budgets—at the cost of maintaining more episodic caches offline. (iv) Figure A1 (f) varies the number of medoid samples in patched prompts, with little effect, suggesting that only a few representatives per cluster are sufficient.
@@ D.1 Efficiency Results
We use a server equipped with 8× NVIDIA A100-40GB (PCIe) GPUs and dual Intel Xeon Platinum 8275CL CPUs, with host-device transfers over PCIe 4.0 x16. Latency is measured on LongMemEval (Wu et al., 2025a) with 100K context length, where each evaluation spans 100 multi-turn queries. We report the average per-turn latency across five runs for both LLaMA3.2-3B and Qwen2.5-3B.
@@ Results.
Table A1 presents the runtime and memory breakdown. Full KV incurs high decoding latency and large peak GPU memory, with LLaMA reaching 68.9ms per turn and 28.4GB memory. Notably, the KV cache storage of LLaMA (11.6GB) even exceeds the model’s own parameter size, underscoring the unsustainable cost of KV caching. By contrast, Qwen2.5-3B exhibits lower KV storage (3.8GB) due to its smaller attention head configuration (Ainslie et al., 2023) (2 heads vs. 8 heads in LLaMA).

### v2

@@ 4.3 Ablation Study
  • Alternative design: replacing episodic cache management with a RAG-like approach that directly inputs clustered conversation segments with the query.
@@ 4.4 Efficiency Analysis
Figure 7(a) reports decoding latency breakdown and peak GPU memory usage, comparing full KV with EpiCache under cache budgets of 2K-8K. Latency is evaluated per turn, consisting of answer decoding [^4] For fair comparison, decoding latency is measured by generating up to 10 tokens per turn. and, for EpiCache, query embedding, centroid matching, and KV cache retrieval. With fewer KV entries stored, EpiCache reduces decoding latency by up to $2.4\times$ and peak GPU memory by $3.5\times$ compared to full KV. These results demonstrate that EpiCache achieves both faster decoding and substantially lower memory usage, since block prefill constrains peak memory to the fixed cache budget during both prefill and decoding.
EpiCache incurs retrieval overhead only when the conversation shifts topics, requiring a new episode cache to be loaded. To measure the frequency of such transitions, we cluster the initial portion of each conversation and embed subsequent turns to track their episode assignments. As shown in Figure 7(b), cumulative transition counts grow sublinearly compared to the Worst Case line (episode switched every turn), indicating that topics often persist across turns. In such cases, the same episodic cache is reused without loading a new one, so no retrieval overhead occurs. As a result, the effective overhead remains under 5% in the per-turn latency breakdown of Figure 7(a), while maintaining robust LongConvQA accuracy. Detailed measurements are provided in Section D.1.
@@
Figure 7: Efficiency Analysis in Multi-Turn Conversation: (a) Per-turn decoding latency and peak GPU memory for full KV (100K) and EpiCache (E=4E{=}4) with LLaMA-3.2-3B. Query Embed and Match: query encoding and centroid matching, KVs Retrieve: loading episodic cache from CPU to GPU memory. (b) Cumulative episode switches in Realtalk with E=4E{=}4, showcasing how often episodes change across multi-turn conversation.
@@ 6 Conclusion
EpiCache is the first framework that combines block-wise prefill with episodic clustering and sensitivity-aware budget allocation to preserve topic-relevant context under a fixed memory budget. Across multiple LongConvQA benchmarks, EpiCache substantially outperforms existing compression methods, demonstrating that efficient multi-turn interaction is feasible even under strict resource constraints and marking a practical step toward memory-efficient conversational AI.
@@ A.1 Dataset
We evaluate EpiCache on three LongConvQA benchmarks: Realtalk (Lee et al., 2025), LoCoMo (Maharana et al., 2024), and LongMemEval (Wu et al., 2025a). Three benchmarks follow the LongConvQA formulation in Section 2.1, where a long conversational history $\mathcal{H}$ is provided and the model is required to answer a sequence of queries $\mathcal{Q}={q_{1},\dots,q_{N_{q}}}$ grounded in dialogue history. This formulation evaluates the answer accuracy of LLMs in a multi-turn conversation.
@@
LoCoMo (Maharana et al., 2024) is a benchmark of long-term conversations, created through a human-machine pipeline where LLM-based agents generate dialogues grounded in distinct personas and temporal event graphs, and human annotators refine them for long-range consistency. The dataset consists of 10 conversations, each spanning up to 35 sessions with around 300 turns. The QA benchmark is divided into five subtasks: (i) Single-hop, (ii) Multi-hop, (iii) Temporal reasoning, (iv) Open-domain knowledge, and (v) Adversarial. Evaluation is conducted with open-ended genetation with F1 score.
We exclude the adversarial subtask for the following reason. This task tests whether a model can recognize unanswerable questions by choosing between a plausible but incorrect answer and a “no such information” response. Under KV cache compression, however, models frequently over-predict the latter, which leads to spuriously high scores. For example, with LLaMA-3.2-3B the adversarial score is only 12.11 under full KV, yet jumps to 49.78 with 4K KVzip compression—an increase that reflects bias rather than genuine improvement.
This behavior contrasts with other subtasks such as temporal or multi-hop reasoning, where compressed caches consistently degrade performance. Because open-source models already struggle on adversarial questions (Maharana et al., 2024), reporting these inflated numbers would give a misleading evaluation of answer quality under compressed KV cache. We therefore omit adversarial results from our main evaluation and defer a more thorough study of unanswerability detection under compression to future work.
@@ A.2 KV Cache Compression Baseline Setup
Following StreamingLLM (Xiao et al., 2024), we retain a fixed number of sink and recent tokens throughout block prefill. Specifically, we fix the number of sink tokens to 128 for all models, while the remaining budget M128M-128 is assigned to the most recent tokens.
@@ A.3 EpiCache Setup
For segment construction, we set the embedding window size $w_{\text{embed}}$ to 4, selected from 2,4,8{2,4,8}. To cluster segments into episodes, we apply the standard K-means algorithm with k-means++ initialization (Arthur & Vassilvitskii, 2007). This offline segmentation and clustering stage completes within one minute, incurring negligible overhead.
For sensitivity-aware budget allocation, we estimate per-layer weights using a single randomly sampled long document from the BookSum (Kryściński et al., 2022) dataset, chosen to avoid bias from any specific conversational dataset. By performing two forward passes—one with the full causal mask and one with the block-prefill mask—we measure layer-wise deviations and compute allocation weights. Because only one sample is used, the overhead of this calibration step is negligible.
@@
For the patched prompt, we use the medoid segment of size 8, selected from 4,8,12{4,8,12}. Ablation studies show that different segment sizes yield only marginal performance differences. After constructing episodic KV caches, we offload them to offline memory (e.g., CPU) to minimize GPU memory usage. During online decoding, episodic caches are retrieved from offline memory. This design enables constrained GPU memory usage while keeping retrieval overhead manageable. A comparison of this design choice is provided in the following section.
@@ Appendix B Alternative Design Exploration
A representative alternative design is a RAG-like approach, where conversation segments are clustered and the query is used to select the most relevant clusters, which are then directly fed into the LLM to form a KV cache under a fixed budget. The key difference from block prefill eviction is that block prefill spans the entire dialogue and selects cluster-related tokens, whereas the RAG-like method isolates only the chosen clusters when constructing the KV cache. As shown in Figure A1 (a), this leads to substantially worse performance—not only below EpiCache but also below block prefill-based compression methods such as KVzip.
This result further suggests that clustering alone is insufficient for directly constructing episodic memory inputs: conversation histories segmented only by clustering remain under-contextualized. Developing more effective clustering or episode partitioning strategies tailored to RAG-like design thus remains an interesting direction for future work.
@@
Figure A1 (b) varies the block size $M_{\text{block}}$ used during block prefill. Accuracy differences remain minor across settings. In practice, smaller block sizes reduce peak memory but require more iterations to span the full context, whereas larger blocks incur higher memory during prefill but process the history more efficiently. This reflects a memory-latency trade-off, and block size can be tuned according to deployment constraints.
@@
Figures Figure A1 (c-f) examine clustering-related hyper-parameters. (i) Figure A1 (c) sweeps the segmentation window size $w_{\text{emb}}$ for utterance grouping, showing little sensitivity and indicating robustness to segmentation granularity. (ii) Figure A1 (d) varies the encoder for clustering and query embedding. The LLM’s own embedding layer yields weaker performance, while lightweight sentence encoders (MiniLM-L6-v2, 0.06B; Qwen3-0.6B) deliver strong gains over InfiniPot. Larger encoders (Qwen3-4B) add only marginal improvements, suggesting that lightweight sentence encoders are sufficient for conversation clustering in EpiCache. (iii) Figure A1 (e) sweeps the number of episodes $E$, where larger $E$ yields finer-grained topic segmentation and more accurate episodic caches—especially under tight budgets—at the cost of maintaining more caches offline. (iv) Figure A1 (f) varies the number of medoid samples in patched prompts, with little effect, suggesting that only a few representatives per cluster are sufficient.
@@ D.1 Efficiency Results
We use NVIDIA A100-40GB (PCIe) GPUs with dual Intel Xeon Platinum 8275CL CPUs. Latency is measured on LongMemEval (Wu et al., 2025a) with 100K context length, where each evaluation spans 100 multi-turn queries. We report the average per-turn latency across five runs for both LLaMA3.2-3B and Qwen2.5-3B.
@@ Results.
Table A1 presents the runtime and memory breakdown. Full KV incurs high decoding latency and large peak GPU memory, with LLaMA reaching 68.9 ms per turn and 28.4 GB memory. The KV cache storage of LLaMA (11.6 GB) also exceeds the model’s parameter size. In comparison, Qwen2.5-3B shows a smaller KV cache size (3.8 GB), which can be attributed to its lower number of KV attention heads (2 vs. 8 in LLaMA) (Ainslie et al., 2023).